# Constructing Benchmarks and Interventions for Combating Hallucinations in LLMs

### v1

@@ Abstract
Large language models (LLMs) are susceptible to hallucination, which sparked a widespread effort to detect and prevent them. Recent work attempts to mitigate hallucinations by intervening in the model’s computation during generation, using different setups and heuristics. Those works lack separation between different hallucination causes. In this work, we first introduce an approach for constructing datasets based on the model knowledge for detection and intervention methods in closed-book and open-book question-answering settings. We then characterize the effect of different choices for intervention, such as the intervened components (MLPs, attention block, residual stream, and specific heads), and how often and how strongly to intervene. We find that intervention success varies depending on the component, with some components being detrimental to language modeling capabilities. Finally, we find that interventions can benefit from pre-hallucination steering direction instead of post-hallucination. The code is available here.
@@ 1 Introduction
A notable issue identified in LLMs is their susceptibility to generate outputs that are not grounded in the input of the model or in real-world facts, or are inconsistent with earlier generations in a session (Ji et al., 2023; Sharma et al., 2023; Kalai & Vempala, 2023). These different issues are typically titled collectively as hallucinations. The importance of addressing these phenomena lies in the reduced reliability of models that sometimes hallucinate. Numerous efforts have been made to tackle these issues, including blackbox approaches that propose suprervised and reinforcement learning finetuning recipes attempting to fix the generated text (Tian et al., 2023; Razumovskaia et al., 2023; Ouyang et al., 2022), as well as whitebox approaches, which aim to detect and prevent hallucinations using the model’s inner states, such as weights and internal activations (Burns et al., 2022; He et al., 2023; Rateike et al., 2023; Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023).
In this work we focus on whitebox approaches, which may enable earlier actions upon hallucination during response generation and propose a better understanding of the hallucination mechanism within LLMs. We claim that a proper benchmark for the task of whitebox intervention should only include hallucinations for cases in which the model “knows” the correct response but still generates it incorrectly, and not cases in which, for example, the model was not exposed to the ground truth knowledge at pretraining, and therefore has no chance to respond correctly. To this end, we first define the different hallucination types w.r.t effective whitebox mitigation. We then propose a method to automatically differentiate between these types and with it construct model-specific benchmarks (each benchmark is tailored to a given model’s internal knowledge). We construct such benchmarks in two settings, one that requires the model to rely on knowledge from the input context within the prompt (termed open-book hallucination) and another that requires the model to rely only on its parametric knowledge (termed closed-book hallucination).
@@
The scenario of closed-book hallucination, wherein the model hallucinates despite knowing the correct answer and lacking any alternative answer in the context, is highly intriguing. Intuitively, one might assume that such hallucinations should not occur. However, examples in the literature demonstrate how user input can trigger such behavior (see Nardo (2023); Wolf et al. (2023)). In Section 3.3, we propose a novel setting to investigate this phenomenon further.
@@
Armed with tailored benchmarks, we turn to whitebox analysis and utilize steering vectors (TurnTrout et al., 2023; Subramani et al., 2022) to mitigate hallucinations. Different from prior work, which detected and created a steering vector for hallucinations after they were generated (Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024), we target detection before hallucination onset, with the goal of steering the generation away from hallucination before it is materialized in surface tokens. In addition, despite some success in detection and mitigation in prior work, there are no clear guidelines on how to do it. We use our framework to create such guidelines by rigorously investigating different choices for intervention such as on which component in the model’s architecture to apply them. We introduce dynamic intervention, performing it only when needed, and compare it to static intervention. As well as using pre-hallucination inner activations for the steering vector instead of post-hallucination. We discover several insights: (i) classification accuracy is not a good approximation to generation accuracy and perplexity should be measured as well; (ii) intervening on the attention component is preferred over the MLP part; (iii) intervention on the residual has a perplexity compromise, which is mitigated to some extent when intervening dynamically;(iv) intervening using pre-hallucination inner states for the steering vector is better than post-hallucination;
@@
Our main contributions are: (I) A methodology for constructing per-LLM benchmarks containing grounded and hallucinated outputs in cases where the model knows the information for a correct response in its parameters, in both open-book and closed-book settings. We will release the datasets for our tested models (Llama2-7b and Goat-7B); (II) A comprehensive investigation into optimal intervention strategies within the model architecture. and a novel dynamic intervention approach that can help intervene only when needed. (III) Insights into the evaluation of interventions.
@@ 2 Problem Setup
Consider a question-answering (QA) setup, where a model is given a question and generates an answer. We consider two settings: open-book QA where the gold-truth answer appears in a given context and closed-book QA where the gold-truth answer resides in the model’s parameters. A common recipe for mitigating hallucinations involves the following steps: (1) constructing a dataset with hallucination labels; (2) training a detector (classifier) on intermediate representations to identify cases of hallucinations; and (3) mitigation - intervening in the model’s computation to prevent a hallucinated answer. The third step requires the detector from step two to decide where to intervene. In this work we focus on steps (1) and (3). See Appendix D for details on how we perform step (2), training the detector.
@@ 2.1 Mitigation Setup
We aim to mitigate hallucinations by adjusting the model’s activations. Following prior work in this area (Li et al., 2023; Hoscilowicz et al., 2024; Chen et al., 2024; Marks & Tegmark, 2023; Zou et al., 2023) and in other tasks (Bau et al., 2019; Ravfogel et al., 2021), our intervention approach is to add a vector to the model’s activations , at some layer and component , where . For the MLP, attention block, and residual, our modification is at the output of the layer, and for the heads, it is at the output of the layer before applying the projection matrix back to the embedding space. The vector is meant to define a linear separation between hallucination and grounded subspaces. Then, during inference, our intervention formula, , shifts the activation vector in the direction of the grounded subspace, hopefully making the model generate a grounded answer. Here emphasizes the impact of the intervention. is sometimes known as a steering vector (Subramani et al., 2022; TurnTrout et al., 2023). We note that steering vectors do not introduce supplementary knowledge to the model, and computing them typically relies on labeled data. This information will be used in deciding how to construct the dataset.
Our paper explores the gory details of the mitigation step, how to compute the steering vector, evaluate its performance, and suggest insights for efficient mitigation.
@@ 2.2 Knowledge Categorization
In closed-book setting Hallucinations can be caused by different reasons, depending on whether or not the model knows the golden response, i.e., the model is expected to induce the golden response tokens from its activations with high confidence. We define three types of knowledge categories that in each hallucination can occur:
1.
The model does not know the golden answer at all.
2.
The model considers the golden answer as one of a few possible ones, without a clear preference for the golden answer.
3.
The model knows the golden answer.
@@
We argue that if the model does not know the golden answer (type 1), it is not possible to prevent hallucination without using external knowledge. If the model has not clear preference for the golden answer (type 2), intervention without external knowledge might be possible, but seems unlikely. Instead, the model could choose to refrain from answering in cases of uncertainty in the answer. This is a different type of intervention that we defer for future work. On the other hand, if the model does know the golden answer, it may be possible to direct its output towards the golden answer via whitebox intervention. Therefore, we focus on type 3 in this work. The automatic way we differentiate between those types of knowledge is explained in Section 3.
@@
This categorization into hallucination from the model knowledge that we propose was not done in prior work on mitigation (Zhang et al., 2024; Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024), and therefore prior benchmarks mixed the 3 types.111One exception is a concurrent work byYuan et al. (2024), which tries to consider the model’s knowledge in creating a dataset. We advocate for the significance of this categorization to address the diverse causes of hallucination effectively.
3 Intervention Dataset Construction
@@
Training and evaluating detectors and steering vectors for mitigating hallucinations necessitates labeled datasets, containing hallucinated and grounded examples. We specifically aim at constructing such datasets for type 3 hallucinations (Section 2.2). To this end, we first describe an automatic method to differentiate between the three hallucination types. Then, we utilize this method to create training, validation, and test sets in open-book and closed-book QA settings for type 3 cases. Figure 1 outlines this process.
3.1 Hallucination Example Selection
@@
To automatically differentiate between the three types of knowledge situations that can create hallucinations, we use repeated sampling from a tested LLM . Concretely, given a QA example consisting of a question and a gold answer, we sample answers from using temperature sampling, and categorize the example according to the number of times the gold response was sampled: 0 times refers to type-1, 1–3 times refers to type-2, and 4–5 times refers to type-3. We retain only type-3 examples, thus tailoring a dataset per model . Here and onwards in the paper, we will use the terminology for type-3 examples as the model knows the answer. From those examples, we will next create a labeled hallucination dataset. See Appendix A.1 for specifics.
This method is easy to implement and we found it useful for splitting into our three types.222One may also train a dedicated model to separate examples by knowledge of a model (Kadavath et al., 2022) , but that is more complicated than our approach. We applied it in both open- and closed-book settings; see below.333This is primarily needed in the closed-book setting but to maintain consistency we applied it in both settings.
@@ 3.2 Open-Book Dataset Creation
In this scenario, we study hallucinations that occur when there is a conflict between a model’s parametric knowledge and a given context. A hallucination in this case is when the model’s answer deviates from information given in the prompt context. We use the DisentQA dataset Neeman et al. (2022), where each example is a question , a text context , and two answers: a parametric answer and a contextual answer . See Figure 2 for an example.
@@
Next, given a full DisentQA example {}, containing a question and context inputs, and the parametric and contextual gold answers, we measure how much the model prefers over : . Where represents the rank assigned by the model to token compared to all other tokens after projection to the vocabulary space at the last layer; a low-rank number indicates a higher preference for answer . If for some threshold then the example is labeled ‘grounded’, and if it is labeled ‘hallucination’. The result of this process is a model-specific open-book DisentQA labeled for hallucinations. See full statistics and more details in Appendices A.1, A.2 and B.
@@ 3.3 Closed-Book Dataset Creation
In this setting, we study hallucinations that occur when an LLM generates an answer that deviates from the world knowledge it has in its parameters.444Note that type 2 could be problematic to use as the model does not clearly prefer the golden answer thus the steering vector may not steer to the golden answer. To this end, we follow prior work that shows that models may simulate different personas, making them hallucinate even though they know the answer (Wolf et al., 2023; Joshi et al., 2023; Janus, 2022). This kind of behavior may happen when a user writes something untrue or misleading (Nardo, 2023; Wolf et al., 2023; Zhang et al., 2023; Xu et al., 2023).
Concretely, we constructed 20 bad-shot examples using ChatGPT (Ouyang et al., 2022). These are false QA pairs, where the false answer is semantically similar to the gold one. For instance, here is a good-shot example and its corresponding bad-shot example:
@@
Three good-shots (or three bad-shots) question: Which colour lies between red and yellow in the visible spectrum? (wrong) answer: Golden answer (): Orange
Then, for a given LLM , we used to answer the questions in TriviaQA (Joshi et al., 2017) in a closed-book setup. As part of the question instruction, we provide the good-shots or bad-shots as few-shot exemplars for the task (Figure 3). For a given example {, } of a question and a gold answer , we label it as ‘grounded’ or ‘hallucination’ via the difference in rank of following the bad-shot or good-shot prompts: . If the rank does not change ( ), the example is considered grounded. If the worsens ( ) and the model predicted the answer correctly given the good-shot, we consider the example as a hallucination. Otherwise, the example is removed. This labeling procedure is illustrated in Figure 4. More details, as well as examples of generated hallucinations using bad shots, are in Appendices A.1, A.3 and B.
@@ 4 Intervention Analysis Approach
With our datasets labeled with hallucinations of type 3, where the model knows the answer but still makes a mistake, we perform a rigorous analysis of whitebox model intervention using steering vectors (See Section 2.1). To this end, we discuss the design choices of intervention via steering vectors in this setting, and reassess proper evaluation metrics, indicating that both classification and generation accuracy are important to measure.
@@ 4.1 Exploring Methodological Variability in Intervention
Various studies employed different settings for intervention, and it remains unclear which approaches are most effective. We conduct experiments with the common choices that appear in most works to establish guidelines, as well as introduce novel settings. Figure 5 presents the variables we explore. Specifically, we identify the following degrees of freedom from the body of work in designing an intervention experiment.
First, a steering vector may intervene in different components: MLP (Zhang et al., 2024), attention block (Zhang et al., 2024), heads (Chen et al., 2024; Li et al., 2023; Hoscilowicz et al., 2024), or residuals (Marks & Tegmark, 2023; Zou et al., 2023)555Zou et al. (2023) also used the attention for a different intervention using LoRRA.. In this work, we will intervene in all the components.
Second, the steering vector is commonly computed using examples containing the generated answer (Li et al., 2023; Hoscilowicz et al., 2024; Chen et al., 2024; Marks & Tegmark, 2023; Zhang et al., 2024; He et al., 2023; Zou et al., 2023). However, it is then used for mitigation before the answer is generated. Thus, there may be a discrepancy between the way it was computed and how it is used. In our experiments, we consider both the common post-answer method and a new pre-answer setup where the steering vector is calculated before answer generation instead of after. Pre-answer is a novel setting we introduce, which is more aligned with how is used during intervention (before hallucination is generated).
Third, the intervention (-compute) may be computed from a detection classifier or from another summary statistic of hallucination vs. grounded cases. In particular, it is common to take the linear separator for the former (Marks & Tegmark, 2023; Chen et al., 2024; Li et al., 2023) and the difference in mean vectors of grounded and hallucination examples for the latter (Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024).666One exception is Zhang et al. (2024), who trained an auto-encoder to learn . In both cases, the decision of where to intervene is made using the detector classifier.
In addition, we introduce another degree of freedom. In static intervention occurs in the same components and layers, for all examples, usually by picking the places that got the highest detection score. This is a common strategy (Li et al., 2023; Marks & Tegmark, 2023; Zhang et al., 2024; He et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024). However, such drastic intervention may not always be necessary. We suggest a different option: dynamic intervention, in which the components to intervene are chosen in each example differently. Our intuition is that intervention should be applied when hallucination is likely to occur, which is pointed at by our detector. Thus, given a threshold , we intervene only in components whose detector accuracy is above (a high-quality detector) and the detector classified the example as hallucination for them.

### v2

@@ Abstract
Large language models (LLMs) are prone to hallucinations, which sparked a widespread effort to detect and prevent them. Recent work attempts to mitigate hallucinations by intervening in the model’s generation, typically computing representative vectors of hallucinations vs. grounded generations, for steering the model’s hidden states away from a hallucinatory state. However, common studies employ different setups and do not properly separate different possible causes of hallucinations, making interventions misguided. In this work, we introduce a method for categorizing examples based on the model’s prior knowledge, named WACK. We construct WACK benchmarks that support interventions in two settings: open-book and closed-book question answering. Using the benchmarks, we perform an extensive investigation of the effect of different choices for intervention, such as the intervened components, and how often and how strongly to intervene. We find that intervention success varies depending on the component, with the attention blocks performing well and the residual stream proving detrimental to language modeling capabilities. We also show that interventions can benefit from representative vectors collected before, rather than after, a hallucination occurs. Finally, we introduce a new dynamic intervention, which intervenes only if needed, and thus is more robust than standard static interventions.111Code and datasets at https://github.com/technion-cs-nlp/hallucination-mitigation.
@@ 1 Introduction
A notable issue identified in LLMs is their prone to generate outputs that are not grounded in the input of the model or in real-world facts, or outputs that are inconsistent with earlier generations in a session [15, 33, 19]. These different issues are typically titled collectively as hallucinations, which are important to address as they reduce the reliability of LLMs. Numerous efforts have been made to tackle these issues, including blackbox methods, which fix the generated text by supervised and reinforcement learning finetuning recipes [38, 32, 28], alongside whitebox methods, which aim to detect and prevent hallucinations using directly the model’s inner states [4, 11, 30, 22, 24, 51].
In this work, we focus on the whitebox approach, which (a) may enable earlier actions upon hallucination during generation; and (b) provides a better understanding of the hallucination mechanism in LLMs. We first categorize the following LLM knowledge types given a question: (1) the model does not know the correct answer (2) the model considers a few options among the correct answer (3) the model knows the correct answer. We claim that a proper benchmark for this task should only include cases of type 3, as hallucinations in this knowledge type could potentially be mitigated without external knowledge. We then propose a method to automatically differentiate between these types for a given model, constructing model-specific benchmarks, where a model has a Wrong Answer despite having Correct Knowledge (WACK). The automation is done by repeat sampling from the model. The final dataset uses only type 3 and labels automatically hallucination/grounded based on the model’s preference.
@@
We present WACK hallucination benchmarks in two settings, one that requires the model to rely on knowledge from the input context in the prompt and not its parametric knowledge (open-book) and one that requires the model to rely only on its parametric knowledge (closed-book). The closed-book setting is intriguing, wherein the model hallucinates despite knowing the correct answer and lacking any alternative answer in the context. Intuitively, one might assume that this should not occur. However, recent work shows how user input can trigger such behavior [26, 42]. We propose a novel setting to examine this by adding irrelevant text with minor mistakes at the start of the prompt, thus triggering a behavior similar to the one a user may create.
@@
Armed with tailored benchmarks, we turn to analysis and utilize steering vectors [40, 37] to mitigate hallucinations, by adding a vector to a hidden state to steer away from hallucinations. We target detection before hallucination onset, intending to steer the generation away from hallucination before it is materialized in surface tokens. Unlike prior work, which detected and created a steering vector for hallucinations after they were generated [22, 24, 51, 7, 12]. Despite some success in mitigation in prior work, there are no clear guidelines on how to intervene. We use our framework and benchmarks to create such guidelines by rigorously investigating different choices for intervention such as which component in the model’s architecture to apply it (MLPs, attention blocks, residual stream, and specific heads). We introduce dynamic intervention, intervening only when needed, and compare it to standard static intervention. We discover several insights: (i) intervening using pre-hallucination inner states for the steering vector is better than post-hallucination; (ii) intervention on the residual has a model’s performance compromise, which is mitigated to some extent when intervening dynamically; (iii) intervening on the attention component is preferred over the MLP part; (iv) model preference between answers (hallucination vs. grounded), which is commonly used as an evaluation metric, is not a good approximation for whether the model will generate the grounded answer;
@@
Our main contributions are: (I) A methodology (WACK) for constructing per-LLM benchmarks containing grounded and hallucination outputs in cases where the model knows the information for a correct response in its parameters, in both open-book and closed-book settings. We released the datasets for our tested models (Llama2-7B and Goat-7B); (II) A comprehensive investigation into optimal intervention strategies within the model architecture; (III) A novel dynamic intervention approach that intervenes only when needed, thus providing a robust solution.
@@ 2 Problem setup
Consider a question-answering (QA) setup, where a model is given a question and generates an answer. We consider two settings: open-book QA where the gold-truth answer appears in a given context and closed-book QA where the gold-truth answer resides in the model’s parameters. A common recipe for mitigating hallucinations involves the following steps: (1) constructing a dataset with hallucination labels; (2) training a detector (classifier) on intermediate representations to identify cases of hallucinations; and (3) mitigation - intervening in the model’s computation to prevent a hallucinate answer. The third step requires the detector from step two to decide where to intervene. In this work we focus on steps (1) and (3). See Appendix D for details on how we perform step (2), as well as various detection results.
@@ 2.1 Mitigation setup
We aim to mitigate hallucinations by adjusting the model’s activations. Following prior work in this area [22, 12, 7, 24, 51] and in other tasks [2, 31], our intervention approach is to add a vector to the model’s activations , at some layer and component , where . For the MLP, attention block, and residual, our modification is at the output of the layer, and for the heads, it is done before applying the projection matrix back to the embedding space. The vector defines a linear separation between hallucination and grounded subspaces. Then, during inference, our intervention formula, , shifts the activation vector in the direction of the grounded subspace, aiming at making the model generate a grounded answer. Here emphasizes the impact of the intervention. is sometimes known as a steering vector [37, 40]. Our work explores the gory details of the mitigation step, how to compute the steering vector, evaluate its performance, and suggest insights for efficient mitigation. Notably, steering vectors do not introduce supplementary knowledge to the model, and computing them typically relies on labeled data. We use this to decide how to construct the dataset.
@@ 2.2 Knowledge categorization
In closed-book setting, Hallucinations can be caused by different reasons, depending on whether or not the model knows the golden response, i.e., the model is expected to induce the golden response tokens from its activations with high confidence. We define three types of knowledge categories that in each hallucination can occur: (1) The model does not know the golden answer at all. (2) The model considers the golden answer as one of a few possible ones, without a clear preference for the golden answer. (3) The model knows and has a clear preference for the golden answer.
@@
We argue that if the model does not know the golden answer (type 1), it is possible to prevent hallucination by using external knowledge or additional training. If the model has no clear preference for the golden answer (type 2), intervention without external knowledge might be possible but seems unlikely. Instead, the model could refrain from answering in cases of uncertainty in the answer. This is a different type of intervention that we defer for future work. On the other hand, if the model does know the golden answer, it may be possible to direct its output toward the golden answer via whitebox intervention, without additional knowledge. Therefore, we focus on type 3 in this work. The automatic way we differentiate between those types of knowledge is explained in Section 3.
@@
This is a novel categorization of hallucination based on knowledge that was not done in prior work on mitigation [50, 22, 24, 51, 7, 12], and therefore prior benchmarks mixed all types.222An exception is a concurrent work [47], which accounted for the model’s knowledge in dataset creation. We advocate for the significance of this categorization to address the diverse causes of hallucination effectively.
3 WACK method for dataset construction
@@
Training and evaluating detectors and steering vectors for mitigating hallucinations necessitates labeled datasets, containing hallucination and grounded examples. We aim to construct such datasets for type 3 knowledge (Section 2.2). To this end, we describe an automatic method to differentiate between the three knowledge types. Then, we create a labeled hallucination dataset in open-book and closed-book QA settings. Figure 1 outlines this process.
3.1 Knowledge example selection
@@
To automatically differentiate between the three types of knowledge situations that can create hallucinations, we use repeated sampling from a tested LLM . Concretely, given a QA example consisting of a question and a gold answer, we sample answers from using temperature sampling, and categorize the example according to the number of times the gold response was sampled: 0 times refers to type-1, 1–3 times refers to type-2, and 4–5 times refers to type-3. We retain only type-3 examples, thus tailoring a dataset per model . henceforth in the paper, we will use the terminology for type-3 examples as the model knows the answer. From those examples, we will next create a labeled hallucination dataset. This method is easy to implement and we found it useful for splitting into our three types.333One may also train a dedicated model to separate examples by model’s knowledge [18], but that is more complicated than our approach. We applied it in both open-book and closed-book settings; see below. This is primarily needed in the closed-book setting but for consistency, we applied it in both settings. See Appendix A.1 for specifics.
@@ 3.2 Open-book dataset creation
In this scenario, we study hallucinations that occur when there is a conflict between a model’s parametric knowledge and a given context. A hallucination in this case is when the model’s answer deviates from information given in the prompt context. We use the DisentQA dataset [27], where each example is a question , a text context , and two answers: a parametric answer and a contextual answer . See Figure 2 for an example.
@@
Next, given a full DisentQA example {}, containing a question and context inputs, and the parametric and contextual gold answers, we measure how much the model prefers over : . Where represents the rank assigned by the model to token compared to all other tokens after projection to the vocabulary space at the last layer; a low-rank number indicates a higher preference for answer . If for some threshold then the example is labeled ‘grounded’, and if it is labeled ‘hallucination’. The result of this process is a model-specific open-book dataset labeled for hallucinations. See details in Appendices A.1, A.2, and B.
@@ 3.3 Closed-book dataset creation
In this setting, we study hallucinations that occur when an LLM generates an answer that deviates from its parametric world knowledge. To this end, we follow prior work that shows that models may simulate different personas, making them hallucinate even though they know the answer [42, 17, 14]. This kind of behavior may happen when a user writes something untrue or misleading [26, 42, 49, 44].
Concretely, we constructed 20 bad-shot examples using ChatGPT [28]. These are false QA pairs, where the false answer is semantically similar to the gold one. Thus, simulating minor mistakes that a user may create.444[45] generated hallucination trigger prompts by constructing them to be out-of-distribution or semantically weak. In contrast, we believe that our bad-shots prompts maintain greater coherence and alignment within the distribution. For instance, here is a good-shot example and its corresponding bad-shot example:
@@
few shots: 3 good-shots (or 3 bad-shots) question: Which colour lies between red and yellow in the visible spectrum? (wrong) answer: Golden answer (): Orange
Then, for a given LLM , we used to answer the questions in TriviaQA [16] in a closed-book setup. As part of the question instruction, we provide the 3 good-shots or 3 bad-shots as few-shot exemplars for the task (Figure 3). For a given example {, } of a question and a gold answer , we label it as ‘grounded’ or ‘hallucination’ via the difference in rank of following the bad-shot or good-shot prompts: . If the rank does not change ( ), the example is considered grounded. If the worsens ( ) and the model predicted the answer correctly given the good-shot, we consider the example as a hallucination. Otherwise, the example is removed. This labeling procedure is illustrated in Figure 4. For more details and examples of hallucinations with bad shots, see Appendices A.1, A.3 and B.
@@ 4 Intervention analysis approach
With our WACK datasets, using knowledge of type 3, where hallucinations occur although the model knows, we perform a rigorous analysis of whitebox model intervention via steering vectors (Section 2.1). We discuss the design choices of intervention and reassess proper evaluation metrics.
@@ 4.1 Exploring methodological variability in intervention
Various studies employed different settings for intervention, and it remains unclear which approaches are most effective. We conduct experiments with the common choices that appear in most works to establish guidelines and introduce novel settings. Figure 5 presents the variables we explore. Specifically, we identify the following degrees of freedom from prior work in designing an intervention.
First, a steering vector may intervene in different components: MLPs [50], attention blocks [50], heads [7, 22, 12], or residual stream [24, 51].555[51] also used the attention for a different intervention using LoRA [13]. In this work, we explore all of the components.
Second, the steering vector is commonly computed using examples containing the generated answer [22, 12, 7, 24, 50, 11, 51]. However, it is then used for mitigation before the answer is generated. Thus, there may be a discrepancy between the way it was computed and how it is used. In our experiments, we consider both the common post-answer method and a new pre-answer setup where the steering vector is calculated before answer generation instead of after. Pre-answer is a novel setting we introduce, which is more aligned with how is used during intervention.
Third, the intervention (-compute) may be computed from a detection classifier or another summary statistic of hallucination vs. grounded cases. In particular, it is common to take the linear separator for the former [24, 7, 22] and the difference in mean vectors of grounded and hallucination examples for the latter [22, 24, 51, 7, 12].666One exception is Zhang et al. [50], who trained an auto-encoder to learn . In both cases, the detector makes an intervention decision.
Additionally, we introduce another degree of freedom. Static intervention occurs in the same components and layers, for all examples, usually by picking the places that got the highest detection score. This is a common strategy [22, 24, 50, 11, 7, 12]. However, such drastic intervention may not always be necessary. We suggest a different option: dynamic intervention, in which the components to intervene are chosen in each example differently. Our intuition is that intervention should be applied when hallucination is likely to occur, which is pointed at by our detector. Thus, given a threshold , we intervene only in components whose detector accuracy is above (a high-quality detector) and the detector classified the example as a hallucination for them.